OpenStudy (anonymous):

6/5+i RATIONALIZE THE DEMONIATOR

OpenStudy (earthcitizen):

\[6/(5+i) *((5-i)/(5-i)) = 30-i6/26\]

OpenStudy (hoblos):

multiply it by (5-i)/(5-i) you get 6(5-i)/25+1 = 6(5-i)/26 = 3(5-i)/13